a monopolistic firm sells it product to two types of customers , business customers and other customers . its total cost are given by Tc=20+2Q , where Q denotes total output . the total demand of business customers is given by Q^3 =14-P .The total demand of other customers is given by Q^T=5-0.5P . Assuming that third degree price discrimination between the two groups is possible. Calculate the prices that the firm will charge each group , the amount it will sell to each group , the consumer surplus of each group and the firm's total profit

Then describe informally and provide intuition for how a ban on price discrimination would affect the consumer surplus of business customers , the consumer surplus of other customers and the firm's profit
Answer #1

We have, T= 20 + QQ othes Customers Business - Customers Q = 14 - P p=14-Q mp = 14-20 MC: 2 MR = mc 14-20 = 2 Q = 5-0-5 P P=By a ban on price discrimination, it eleminate the motive of impasation of price dis somation Cie) elemination or reducing co